OpenStudy (anonymous):

f(x) = x3+11x2+39x+29 What would be the zeros for this function?

OpenStudy (anonymous):

if you graph the function, u can determine where it intersects the x-ax.

OpenStudy (ybarrap):

x = -5-2 i x = -5+2 i x = -1